OpenStudy (anonymous):

Describe how to find the coordinates of the midpoint of a segment in 3D space

OpenStudy (zarkon):

average coordinate wise.

OpenStudy (anonymous):

what?

OpenStudy (zarkon):

average coordinate wise the two endpoints fore example if the endpoints are (1,2,3) and (5,4,7) the midpoint is \[\left(\frac{1+5}{2},\frac{2+4}{2},\frac{3+7}{2}\right)\] \[=(3,3,5)\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (zarkon):

np